What is the normal dosage for azithromycin, an antibiotic, is 4.5mg/2 kg of body weight. The patient weighs 150 lbs. How many grams of the medication is needed for one dose?
Answer #1

ANSWER.

* Calculation of grams of medication required for this patient.

         Recommended dosage = 4.5 mg/2kg.

                                                = 2.25 mg/kg.

         Weight of the patient = 150 lbs.

Convert lbs to kg for calculation

               1 lbs = 0.454 kg.

      So

          Weight of the patient = 150 * 0.454

                                            = 68.1 kg

      Dose required for this patient in mg = 2.25 * 68.1

                                                                = 153.23 mg

         Convert mg to gm for calculation

                        1 mg = 0.001 gm.

         So

     Dose required for this patient in gm = 153.23 * 0.001

                                                                           = 0.153 gm